The peso is the currency of Argentina. Suppose that the peso/$ forward exchange rate, F, is exactly equal to traders' expectations for the value of the peso/$ spot rate, E, in six months' time. a. If F = 1.15 and E= 1.10, what is the forward discount on pesos? b. Assuming that interest rate parity (IRP) holds, what would you expect to be true of the interest differential between six-month dollar deposits and six-month peso deposits, that is, R$ - Rpeso? c. Argentina has a long history of defaulting on its foreign debt. Do you think the IRP condition is likely to hold for dollar- versus peso-denominated deposits? If not, is the interest differential R$ - Rpeso likely to be greater or less than what is predicted by IRP?
Answer #1

a. Forward discount (annualized) = (Forward-Spot)/Spot * 360/No.of days *100

Forward = 1.15, spot = 1.10 and No. of days = 180 days (six months)

Annualized forward discount = (1.15-1.10)/1.10*360/180 *100 = 9.09%

Annualized forward discount 9.09%b.

b. According to IRP, six month interest rate differential should be equal to the six month forward difference

Therefore, l = (1.15-1.10)/1.10 = 0.04545 = 4.55%

six month interest rate differential = 4.55%

c. No. IRP is not likely to hold for dollar vs peso denominated deposits. This is because the long default history of Argentina is likely to cause a greater interest rate differential than that indicated bu interest rate parity.